OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x2 − 16x + ____ Numerical Answers Expected!

OpenStudy (anonymous):

what is half of \(16\) ?

OpenStudy (anonymous):

8.

OpenStudy (anonymous):

k what is \(8^2\) ?

OpenStudy (anonymous):

i.e. what is \(8\times 8\) that is your answer

OpenStudy (the_fizicx99):

Hm, Well you know that the middle term is 2(a)(b), or at least I assume you do.. x^2 - 16x + __ √x^2 = x -16x = 2(x)(b) = 2xb -16x = 2xb ( solve for b) /2x /2x -8 = b (x - 8)^2 = (x)^2 - 2(x)(8) - (8)^2 = x^2 - 16x + 64